OpenStudy (anonymous):

Evaluate cos(sin^-1 0)

sam (.sam.):

cos(sin^-1 0) remember, sin^-1 0=0,180,360 so, =cos(sin^-1 0) =cos(0) , cos180, cos360 =1,-1

OpenStudy (anonymous):

so it's 1, -1?

OpenStudy (anonymous):

what is arcsin in terms of sin or cos?

sam (.sam.):

yes

myininaya (myininaya):

arcsin(0) is only equal to 0 since -pi/2<=arcsin(x)<=pi/2

myininaya (myininaya):

so cos(0)=1 like you said above